OpenStudy (anonymous):

Help guys ! :( Ill give a medal to the helper . find the equation of the line indicated point with slope m

OpenStudy (anonymous):

(5,0) , m = - 2/3

OpenStudy (anonymous):

@UmarsBack

OpenStudy (anonymous):

I think i need to use the Point Slope Formula

OpenStudy (anonymous):

@annas mada'ame How did you do that?

OpenStudy (anonymous):

oops forgot to take out value of b

OpenStudy (anonymous):

oh and I'm sir xD

OpenStudy (anonymous):

final equation will be y=-2/3x+10/3

OpenStudy (anonymous):

sir help me please

OpenStudy (anonymous):

I took that out from y = mx+b equation of a line

OpenStudy (anonymous):

To start, you know what m is; it's just the slope, which you said was -2/3. So you can right away fill in the equation for a line somewhat to read: y=-2/3x+b.

OpenStudy (anonymous):

I cant use the equation thingy

OpenStudy (anonymous):

To find b, think about what your (x,y) point means, (5,0) When x of the line is 5, y of the line must be 0. Because you said the line passes through this point right? Now, look at our line's equation so far b is what we want, the -2/3 is already set and x and y are just two "free variables" sitting there. We can plug anything we want in for x and y here, but we want the equation for the line that specfically passes through the the point (5,0) So, why not plug in for x the number 5 and for y the number 0? This will allow us to solve for b for the particular line that passes through the point you gave!. (5,0) y=mx+b or 0=-2/3 × 5+b, or solving for b: b=0-(-2/3)(5) => b=10/3.

OpenStudy (anonymous):

so the final eq you get is y=-2/3x+10/3

OpenStudy (anonymous):

I hope that explains :)

OpenStudy (anonymous):

@Dannica22

OpenStudy (anonymous):

Look at this Given P(-7 , 1) and m = 3/4 (Use Point Slope Formula y-y1=m(x-x1) y-1=3/4(x+7) 4(y-1)=3(x+7) 4y-4=3x+21 so therefore : 3x-4y+25=0

OpenStudy (anonymous):

@annas my bad ! hehehe not reading my notes. your point is I will use the Slope intercept Formula hehehehehe

OpenStudy (anonymous):

:) sure and solution for second question is y=3/4x+25/4

OpenStudy (anonymous):

wait Ill try it my self hehehehe

OpenStudy (anonymous):

try one question by yourself cause as much as you practice your maths gets great :)

OpenStudy (anonymous):

Can you make that equation into this equation x+y=0

OpenStudy (anonymous):

On the second question I use Point slope Formula

OpenStudy (anonymous):

sure y=3/4x+25/4 -3/4x + y = 25/4 -3x + 4y = 25 -(3x - 4y)= 25 3x - 4y = -25 3x - 4y + 25 = 0

OpenStudy (anonymous):

great :) you are pretty smart kido :)

OpenStudy (anonymous):

Yeahhhhh it goes that thing !!!!!!!!!!!!!! Thank you !! But hahaha its kinda confusing a little but it help me a lot . Why is there something wrong with the Equation thingy in here?

OpenStudy (anonymous):

I never tried equation thing xDD & no worries we are come here to learn & help :)

OpenStudy (anonymous):

SO the first question with the answer of y=-2/3x+10/3 .. The equation goes like this 2/3x-y+10/3 = 0

OpenStudy (anonymous):

y=-2/3x+10/3 2/3x + y = 10/3 2x + 3y = 10 2x + 3y - 10 = 0

OpenStudy (anonymous):

Thank you (y)

OpenStudy (anonymous):

anytime :)